OpenStudy (anonymous):

1. Solve, using the substitution method. m + n = 5 m – n = 3 There is no solution. The solution is (2, –1) The solution is (4, 1) There are an infinite number of solutions.

OpenStudy (texaschic101):

m + n = 5 ---> m = 5 - n now sub 5 - n in for m in the other equation m - n = 3 (5 - n) - n = 3 5 - 2n = 3 -2n = 3 - 5 -2n = -2 n = 1 now sub 1 in for n into either of the original equations m + n = 5 m + 1 = 5 m = 5 - 1 m = 4 The solution is (4,1)

OpenStudy (anonymous):

thx a million man :D